# SlumpGuard: An AI-Powered Real-Time System for Automated Concrete Slump Prediction via Video Analysis

### v1

@@ 5.4.1 Model Training Setup
The model was optimized using the AdamW optimizer [19] with a learning rate of $41e-$ 4, and a weight decay of $41e-$ 4. We used a total batch size of 128 and trained for 10 epochs using the OneCycle learning rate scheduler [20] with cosine annealing, with a peak learning rate set to $31e-$ 3.
@@
Table 3: Ablation study results on the training strategy for R(2+1)D-18. ✓indicates that the technique was applied, while a blank cell denotes that it was not.
@@ 6 Conclusion
    To facilitate video-based concrete slump prediction, we constructed a dataset comprising 6,443 video clips, each 10 seconds long, capturing various concrete pouring scenes. Bounding box annotations for the discharge chute of the concrete mixer were provided to enable the development of deep learning-based methods.

### v2

@@ 5.4.1 Model Training Setup
The model was optimized using the AdamW optimizer [39] with a learning rate of 1e41e-4, and a weight decay of 1e41e-4. We used a total batch size of 128 and trained for 10 epochs using the OneCycle learning rate scheduler [40] with cosine annealing, with a peak learning rate set to 1e31e-3.
@@
Table 3: Ablation study results on the training strategy for R(2+1)D-18. “✓” indicates that the technique was applied, while a blank cell denotes that it was not.
@@ 6 Conclusion
    To facilitate video-based concrete slump prediction, we constructed a dataset comprising 6,443 video clips, each 10 seconds long, capturing various concrete pouring scenes. Bounding box annotations for the discharge chute of the mixer truck were provided to enable the development of deep learning-based methods.